OpenStudy (anonymous):

Find the derivative y=(6x-1)(5x-2)^-1

OpenStudy (anonymous):

I used the quotient rule- -7/(6x-1)^2

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

can u show the work

OpenStudy (anonymous):

\[\left(\begin{matrix}(6x-1) \\ (5x-2)\end{matrix}\right)=\left(\begin{matrix}(6)(5x-2)-(6x-1)(5) \\ (5x-2)^{2}\end{matrix}\right)=\left(\begin{matrix}30x-12-30x+5 \\ (5x-2)^{2}\end{matrix}\right)=\left(\begin{matrix}-7 \\(5x-2)^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

i got it wrong its actually -7/(5x-2)^2